OpenStudy (anonymous):

Solve. 0.9x +2 <= 1.3 - 3 Solution is an ordered pair.

OpenStudy (saifoo.khan):

\[\Huge \begin{array}l\color{red}{\text{H}}\color{orange}{\text{E}}\color{#9c9a2e}{\text{Y}}\color{green}{\text{Y}}\color{blue}{\text{Y}}\color{purple}{\text{Y}}\color{purple}{\text{Y}}\color{red}{\text{Y}}\color{orange}{\text{Y}}\color{#9c9a2e}{\text{Y}}\color{green}{\text{Y}}\color{blue}{\text{Y}}\color{purple}{\text{Y}}\color{purple}{\text{Y}}\color{red}{\text{Y}}\color{orange}{\text{Y}}\color{#9c9a2e}{\text{!}}\color{green}{\text{}}\end{array}\]

OpenStudy (anonymous):

HIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIII

OpenStudy (anonymous):

Hi JT

OpenStudy (saifoo.khan):

something is missing.

jimthompson5910 (jim_thompson5910):

Hello 0.9x+2<=1.3-3 ... Start with the given inequality. 10(0.9x+2)<=10(1.3-3) ... Multiply both sides by 10 to clear out the decimals. 9x+20<=13-30 ... Distribute and multiply. 9x+20<=-17 ... Combine like terms on the right side. 9x<=-17-20 ... Subtract 20 from both sides. 9x<=-37 ... Combine like terms on the right side. x<=-37/9 ... Divide both sides by 9 to isolate x. So the solution in interval notation is \[\large (-\infty,-\frac{37}{9}]\]

OpenStudy (saifoo.khan):

i think x is missing on the right side.

OpenStudy (anonymous):

0.9x + 2 <= 1.3x - 3

OpenStudy (saifoo.khan):

i knew it.

OpenStudy (saifoo.khan):

LOL.

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

I am not perfect guys.

jimthompson5910 (jim_thompson5910):

0.9x+2<=1.3x-3 ... Start with the given inequality. 10(0.9x+2)<=10(1.3x-3) ... Multiply both sides by 10 to clear out the decimals. 9x+20<=13x-30 ... Distribute and multiply. 9x<=13x-30-20 ... Subtract 20 from both sides. 9x-13x<=-30-20 ... Subtract 13x from both sides. -4x<=-30-20 ... Combine like terms on the left side. -4x<=-50 ... Combine like terms on the right side. x>=(-50)/(-4) ... Divide both sides by -4 to isolate x. note: Remember, the inequality sign flips when we divide both sides by a negative number. x>=25/2 ... Reduce.

jimthompson5910 (jim_thompson5910):

So the answer in interval notation is \[\large [\frac{25}{2},\infty)\]

OpenStudy (anonymous):

This is and ordered pair.

OpenStudy (saifoo.khan):

He got that right!

OpenStudy (anonymous):

What are those little tiny o's?

jimthompson5910 (jim_thompson5910):

that's the infinity symbol

jimthompson5910 (jim_thompson5910):

are you sure that an ordered pair is expected? Or might there be another typo somewhere?

OpenStudy (saifoo.khan):

LOL.

OpenStudy (anonymous):

The tiny oo's are not an option

OpenStudy (saifoo.khan):

\[\Huge \infty\]

OpenStudy (anonymous):

Let me do this again.

OpenStudy (anonymous):

0.9x + 2 <= 1.3 - 3

OpenStudy (anonymous):

Hello?

OpenStudy (saifoo.khan):

is that right?

OpenStudy (anonymous):

Jim? He only confused me/

OpenStudy (anonymous):

Oh, I have to go take a brake. brb

jimthompson5910 (jim_thompson5910):

sry about that

jimthompson5910 (jim_thompson5910):

usually when they want "ordered pairs" as answers, it's for systems and not single inequalities like this

jimthompson5910 (jim_thompson5910):

If you can, post a screenshot of the exact problem so we can see exactly what it says.

OpenStudy (anonymous):

I am using 2 computers

OpenStudy (saifoo.khan):

o.O

OpenStudy (anonymous):

a pc and a macbook

OpenStudy (saifoo.khan):

sooo.. where are u stuck?

OpenStudy (anonymous):

sk....I need to talk to you privately

OpenStudy (saifoo.khan):

There?